# CoMPaSS: Enhancing Spatial Understanding in Text-to-Image Diffusion Models

### v1

@@ 3.2 Token Encoding Ordering Module
Table 1: Most similar prompt variation retrieved by different text encoders.
Text Encoder Used by $F_{\text{text}}$ Space Most similar prompt variation Correct
rephrased negated relation swapped entities
CLIP ViT-L [radford_learning_2021] All 768 1 5088 1231 0.02%
OpenCLIP ViT-H [cherti_reproducible_2023] SD2.1 1024 0 6054 266 0%
OpenCLIP ViT-bigG [cherti_reproducible_2023] SDXL [podell_sdxl_2023], SD3 [esser_scaling_2024], SD3.5 1280 2 6067 251 0.03%
T5-XXL [raffel_exploring_2020] SD3 [esser_scaling_2024], SD3.5, FLUX.1 4096 306 4777 1237 4.84%
@@
Text-to-image diffusion models rely on text encoders to transform natural language descriptions into semantic representations that guide the image generation process. For generation of accurate spatial relationships, these semantic representations must preserve the accurate spatial relationships described in the input text. This is the cornerstone for effective spatial understanding of T2I diffusion models.
Analysis of Spatial Understanding.
@@
We investigate whether the spatial information is properly represented for the diffusion model to understand it. Our key insight is that if text encoders were able to properly preserve spatial relationships, then logically equivalent spatial descriptions should yield high similarities in their encoded representations. A proxy task is designed to quantify this property: given a base prompt describing a spatial relationship (e.g., “A to the left of B”), we generate three variations of it: (1) rephrased: logically equivalent but differently worded (“B to the right of A”), (2) negated relation: substituting the relation phrase with its opposite (“A to the right of B”), and (3) swapped entities: exchanging object positions (“B to the left of A”). Theoretically, only the rephrased variation should produce an encoding similar to the base prompt.
@@
To ensure comprehensive evaluation, we conduct this analysis using all 80 object categories from COCO [lin_microsoft_2014], with four fundamental spatial relationships (“left”, “right”, “above”, “below”), yielding 6,320 prompts. We retrieve the most similar prompt variation according to the similarity between the encoded representations across four text encoders. The results are shown in Tab. 1: even T5-XXL [raffel_exploring_2020] with 11B parameters fails to recognize the logically equivalent variation over 95% of the time. Lighter encoders like CLIP [radford_learning_2021, cherti_reproducible_2023] show near-complete failures across different model sizes, suggesting that current text encoders do not sufficiently preserve spatial relationships in their encoded representations.
Solution.
Figure 4: Overview of Token Encoding Ordering (TENOR). TENOR injects token ordering information into every text-image attention operation in either UNet- or MMDiT-based diffusion models.
@@
Based on this analysis, we propose TENOR, a plug-and-play module designed to compensate for the limitations of current text encoders. TENOR operates by augmenting the conditioning text signals with their original token ordering information in attention operations throughout the text-to-image generation process. Unlike the original positional encoding in transformers which only adds positional encodings to the initial token embeddings [vaswani_attention_2017], TENOR explicitly injects the token ordering information into every text-image attention operation in the diffusion model. This design ensures that spatial semantics are actively preserved every time text guidance influences the image generation process.
Implementation and Efficiency.
@@
Fig. 4 illustrates the overall process of TENOR. Specifically, in every text-image attention operation in the diffusion model, TENOR adds absolute positional encodings [vaswani_attention_2017] to the $K$ vector for UNet-based models [rombach_high-resolution_2022, podell_sdxl_2023], and both $K_{\text{text}}$ and $Q_{\text{text}}$ for MMDiT-based models [peebles_scalable_2023, esser_scaling_2024]. While concurrent approaches have suggested fine-tuning text encoders [chatterjee_getting_2024] during training, performing test-time optimizations [chefer_attend-and-excite_2023, chen_training-free_2023], or incorporating large language models [hu_ella_2024, wu_paragraph--image_2023], such methods add significant computational overhead either during training or at inference time. In contrast, our solution requires no additional trainable parameters beyond the diffusion model itself, and inflicts practically zero computational penalty at inference time. While adapting TENOR requires a brief fine-tuning phase of the diffusion model, this process is significantly more efficient than training new spatial-aware text encoders. Empirically, we find that this approach substantially improves spatial understanding of both UNet-based models and the MMDiT-based FLUX.1.
@@
Figure 5: Qualitative results of models enhanced with CoMPaSS. CoMPaSS enhances the spatial understanding of diffusion models across UNet-based models (SD1.4, SD1.5, SD2.1) and the MMDiT-based FLUX.1.
@@ 4 Experiments
4.1 Training Details
We validate the effectiveness of CoMPaSS on four popular open-weight text-to-image diffusion models, where SD1.4, SD1.5, and SD2.1 are based on the UNet architecture [rombach_high-resolution_2022], and FLUX.1 is based on the MMDiT architecture [peebles_scalable_2023, esser_scaling_2024]. For UNet-based diffusion models, we train the entire UNet with a learning rate of 5e-6. For MMDiT-based FLUX.1, we train a rank-16 LoRA [hu_lora_2021] with a learning rate of 1e-5 due to limited computational resources. All experiments use an effective batch size of 4. Detailed experimental settings can be found in Appendix A.
4.2 Evaluation Metrics
Table 2: Comparing models enhanced with CoMPaSS and their base models on various spatial-related benchmarks. CoMPaSS sets new state-of-the-arts with substantial gains on every spatial-related benchmark, including VISOR [gokhale_benchmarking_2023], T2I-CompBench Spatial [huang_t2i-compbench_2023-1] (T. Spatial), and Geneval Position [ghosh_geneval_2023] (G. Position).
Method VISOR (%) T. Spatial G. Position
uncond cond 1 2 3 4
SD1.4 18.81 62.98 46.60 20.11 6.89 1.63 0.12 0.03
SD1.4 +CoMPaSS 57.41 87.58 83.23 67.53 49.99 28.91 0.34 0.46
SD1.5 17.58 61.08 43.65 18.62 6.49 1.57 0.08 0.04
SD1.5 +CoMPaSS 61.46 93.43 86.55 72.13 54.64 32.54 0.35 0.54
SD2.1 30.25 63.24 64.42 35.74 16.13 4.70 0.13 0.07
SD2.1 +SPRIGHT [chatterjee_getting_2024] 43.23 71.24 71.78 51.88 33.09 16.15 0.21 0.11
SD2.1 +CoMPaSS 62.06 90.96 85.02 71.29 56.03 35.90 0.32 0.51
FLUX.1 37.96 66.81 64.00 44.18 28.66 14.98 0.18 0.26
FLUX.1 +CoMPaSS 75.17 93.22 91.73 83.31 72.21 53.41 0.30 0.60
Table 3: Comparison of image fidelity metrics on FID [heusel_gans_2017] and CMMD [jayasumana_rethinking_2024].
@@ 4.2 Evaluation Metrics
SD1.4 13.20 0.5618
SD1.4 +CoMPaSS 10.67 0.3313
SD1.5 12.82 0.5548
SD1.5 +CoMPaSS 10.89 0.3235
SD2.1 21.65 0.6472
SD2.1 +SPRIGHT [chatterjee_getting_2024] 16.15 0.512
SD2.1 +CoMPaSS 16.96 0.4083
FLUX.1 27.96 0.8737
FLUX.1 +CoMPaSS 26.40 0.6859
We extensively evaluate CoMPaSS on several well-known benchmarks, which are described below:
@@
  • T2I-CompBench [huang_t2i-compbench_2023-1] and GenEval [ghosh_geneval_2023] are two benchmarks designed for comprehensive evaluation of the text-image alignment capabilities of T2I models. In addition to spatial-related evaluations, they provide methodologies for assessing correctness of attribute binding, non-spatial relationships, and counting. We use them to evaluate the performance impact of CoMPaSS on tasks other than spatial understanding.
  • Fréchet Inception Distance (FID) [heusel_gans_2017] and CLIP Maximum Mean Discrepancy (CMMD)  [jayasumana_rethinking_2024] estimate the fidelity of images generated from T2I models. FID estimates the photorealism of generated images. CMMD is an estimator of image quality that aligns better with human preferences and aims to account for some of the limitations of FID. We report them both.
@@ 4.3 Main Results
Spatial Understanding.
We evaluate spatial relationship generation accuracy across multiple benchmarks and report them in Tab. 2. Adding CoMPaSS to existing diffusion models sets new state-of-the-art across all spatial-related metrics. Notably, with the best open-weight diffusion model FLUX.1, CoMPaSS obtains substantial relative gains on VISOR (+98%), T2I-CompBench Spatial (+67%), and GenEval Position (131%). As evident in Fig. 5, while existing models often struggle with spatial relationships, CoMPaSS effectively addresses these limitations even on novel spatial configurations unseen during training, demonstrating significant boosts in spatial understanding.
General Generation Capability and Fidelity.
Apart from spatial-related metrics, we evaluate other tasks in the GenEval [ghosh_geneval_2023] and T2I-CompBench [huang_t2i-compbench_2023-1] benchmarks, and image fidelity scores on FID [heusel_gans_2017] and CMMD [jayasumana_rethinking_2024]. The results are reported in Tabs. 5 and 3. CoMPaSS exclusively improves spatial-related performance, while improving the overall alignment scores and image fidelity. We conjecture that in the base models, the spatial terms were entangled with unrelated semantics due to severely flawed data and unclear representation. By exclusively disentangling the spatial terms, the model also learns to better understand other aspects of language, resulting in improvements on other tasks.
@@ 4.4 Ablation Studies
Table 4: Ablation studies of each component of CoMPaSS. (i) original models; (ii) trained with the SCOP dataset described in Sec. 3.1; (iii) our full method. T2I-CompBench Spatial (T. Spatial) and GenEval Position (G. Pos) scores are reported.
@@
Table 5: Evaluation results of general generation capabilities across a wide range of tasks on GenEval [ghosh_geneval_2023] and T2I-CompBench [huang_t2i-compbench_2023-1]. CoMPaSS exclusively improves spatial-related performance, while improving the overall alignment scores.
Method T2I-CompBench GenEval
Color Shape Texture Spatial Non-Spatial Position Single O. Two O. Counting Colors Attr. Overall
SD1.4 0.38 0.36 0.42 0.12 0.31 0.03 0.98 0.41 0.34 0.74 0.06 0.43
SD1.4 +CoMPaSS 0.49 0.43 0.53 0.34 0.31 0.46 0.99 0.68 0.34 0.73 0.17 0.56
SD1.5 0.38 0.37 0.42 0.08 0.31 0.04 0.96 0.38 0.36 0.75 0.06 0.42
SD1.5 +CoMPaSS 0.50 0.43 0.52 0.35 0.31 0.54 0.99 0.69 0.34 0.72 0.15 0.57
SD2.1 0.51 0.42 0.49 0.13 0.31 0.07 0.98 0.51 0.44 0.85 0.17 0.50
SD2.1 +SPRIGHT [chatterjee_getting_2024] - - - 0.21 - 0.11 0.99 0.59 0.49 0.85 0.15 0.51
SD2.1 +CoMPaSS 0.55 0.43 0.54 0.32 0.30 0.51 0.99 0.69 0.20 0.71 0.15 0.54
FLUX.1 0.69 0.48 0.63 0.18 0.31 0.26 0.92 0.77 0.71 0.66 0.27 0.60
FLUX.1 +CoMPaSS 0.83 0.59 0.71 0.30 0.32 0.60 0.99 0.87 0.71 0.80 0.76 0.76
Figure 6: Ablation study on the TENOR module. TENOR helps the model to gain better generalization capability beyond training data.
@@
In the following, we conduct controlled experiments on SD1.5 (most popular and accessible UNet-based diffusion model) and FLUX.1-dev (a state-of-the-art diffusion model based on the MMDiT architecture) to better understand how each component contributes to generation performance. We report the spatial accuracy metrics under various settings in Tab. 4.
Effect of SCOP.
@@
Comparing settings (i) and (ii) within each model reveals that SCOP alone contributes substantially to the spatial understanding of diffusion models. This finding is in line with our initial motivation, where we identified that one of the main reasons current models are failing spatial-related generations is that current datasets yield severely flawed spatial-related image-text data. It also confirms that SCOP is able to curate a set of high-quality data that directly contributes to the spatial understanding of diffusion models.
Effect of TENOR.
@@
Tab. 4 shows that our full method (setting (iii) within each model) further improves the spatial accuracy. We prompt the models with text descriptions unseen during training, and compare the generation results with setting (ii). The results are shown in Fig. 6. While SCOP provides rich and accurate spatial priors, it is only after incorporating TENOR that the model can better generalize to hard unseen spatial configurations. This phenomenon connects closely to our initial finding about the spatial understanding deficiency of popular text encoders in Tab. 1: when the semantic representations of completely opposite text descriptions are indistinguishable, the model is trained to align mixed image signals from seemingly similar text conditions, resulting in poor generalization beyond training data. TENOR compensates this issue by actively injecting token ordering information into every text-image attention operation throughout the diffusion model, disentangling different spatial configurations at the input level, alleviating the model’s burden of distinguishing different spatial terms, as a result achieving better zero-shot generation on unseen settings.
5 Conclusion
We present CoMPaSS, a versatile training framework that enhances spatial understanding of any text-to-image diffusion model. CoMPaSS addresses the problem of ambiguous spatial relationships in datasets through the SCOP data engine, which curates spatially-accurate training data through principled constraints. CoMPaSS additionally uses a parameter-free module, TENOR, that further boosts spatial understanding of T2I diffusion models of any architecture while adding negligible computational overhead. We conduct extensive experiments on four popular open-weight diffusion models, covering both UNet-based models (SD1.4, SD1.5, SD2.1) and the MMDiT-based FLUX.1, and set new state-of-the-art across well-known benchmarks with substantial relative gains, including VISOR [gokhale_benchmarking_2023] (+98%), T2I-CompBench Spatial [huang_t2i-compbench_2023-1] (+67%), and GenEval Position [ghosh_geneval_2023] (+131%). Importantly, the enhancement in spatial understanding comes without compromising the general generation capabilities or image quality. By addressing a fundamental limitation of current text-to-image diffusion models, we believe CoMPaSS takes an important step towards reliable spatially-accurate image generation, and opens up new possibilities for applications requiring precise spatial control.

### v2

@@ 3.2 Token Encoding Ordering Module
Text-to-image diffusion models rely on text encoders to transform natural language descriptions into semantic representations that guide the image generation process. To generate accurate spatial relationships, these representations must faithfully preserve the spatial information described in the input text. This capability is a cornerstone for effective spatial understanding in T2I diffusion models.
@@
Analysis of Spatial Understanding. To investigate this potential failure point, we conduct a systematic analysis of how well current text encoders represent spatial information. Our key insight is that if text encoders were able to properly preserve spatial relationships, then logically equivalent spatial descriptions should yield high similarities in their encoded representations. To quantify this property, we design a proxy task: given a base prompt describing a spatial relationship (e.g., “A to the left of B”), we generate three variations of it: (1) rephrased: logically equivalent but differently worded (“B to the right of A”), (2) negated relation: substituting the relation phrase with its opposite (“A to the right of B”), and (3) swapped entities: exchanging object positions (“B to the left of A”). Ideally, an encoder that understands the spatial layout should yield the highest similarity for the rephrased variation.
@@
To ensure comprehensive evaluation, we conduct this analysis using all 80 object categories from COCO [43], with four fundamental spatial relationships (“left”, “right”, “above”, “below”), yielding 6,320 prompts. For each base prompt, we identify which of the three variations yields the most similar embedding, testing this across four different text encoders. The results are shown in Tab. 1: even T5-XXL [56] with 11B parameters fails to recognize the logically equivalent variation over 95% of the time. Lighter encoders like CLIP [55, 14] show near-complete failures across different model sizes, suggesting that current text encoders do not sufficiently preserve spatial relationships in their encoded representations.
Table 1: Most similar prompt variation retrieved by different text encoders. Each encoder is used by at least one popular open-weight diffusion model (SD1.5/2.1/3/XL, FLUX.1).
Text Encoder Most Similar Variation Correct
Rephrased Neg. Rel. Swp.
CLIP ViT-L [55] 1 5088 1231 0.02%
OpenCLIP ViT-H [14] 0 6054 266 0%
OpenCLIP ViT-bigG [14] 2 6067 251 0.03%
T5-XXL [56] 306 4777 1237 4.84%
@@
Solution. Based on this analysis, we propose TENOR, a plug-and-play module designed to address the structural ambiguity limitations of current text encoders. TENOR operates by augmenting the text conditioning signals with their original token ordering information within the text-image attention layers of the diffusion model. Unlike standard positional encodings in transformers, which are added only once to the initial token embeddings [68], TENOR explicitly injects this token ordering information into every text-image attention operation. This design ensures that the prompt’s structural and sequential information is consistently available every time text guidance influences the image generation process.
Figure 4: Overview of Token Encoding Ordering (TENOR). TENOR injects token ordering information into every text-image attention operation in both UNet- and MMDiT-based diffusion models.
@@
Implementation and Efficiency. Fig. 4 illustrates the overall process of TENOR. Specifically, for UNet-based models [58, 53], TENOR adds absolute positional encodings [68] to the key ($K$) vectors in each text-image attention layer. For MMDiT-based models [51, 20], it adds these encodings to both the text query ($Q_{\text{text}}$) and key ($K_{\text{text}}$) vectors. Alternative approaches for improving spatial control involve fine-tuning the text encoder [6], performing test-time optimization [7, 13], or incorporating large language models [33, 74]. However, these methods typically introduce significant computational overhead during either training or inference. In contrast, our solution requires no additional trainable parameters and incurs negligible computational overhead at inference time. While adapting TENOR requires a brief fine-tuning phase of the diffusion model, this process is significantly more efficient than training new spatial-aware text encoders. Empirically, we find that this approach substantially improves spatial understanding of both UNet-based models and the MMDiT-based FLUX.1.
It is crucial to clarify the distinct roles of TENOR and the SCOP dataset. TENOR itself is agnostic to the semantic meaning of spatial terms like “left” or “right”. Its function is to provide a clear, unambiguous structural signal about the token order that is often lost by standard text encoders (as shown in Tab. 1). By explicitly preserving the sequence, TENOR ensures that prompts like “A left of B” and “B left of A” produce structurally different conditioning signals for the diffusion model. During fine-tuning, the model then learns to associate these distinct structural signals with the correct visual outcomes provided by the high-quality, unambiguous image-text pairs from our SCOP dataset. This synergy is what enables the significant improvement in spatial understanding: TENOR providing the structural capacity to differentiate prompts, while SCOP provides the semantic ground truth.
@@
Figure 5: Qualitative results of models enhanced with CoMPaSS. Our method improves the spatial understanding of both UNet-based models (SD1.4, SD1.5, SD2.1) and the MMDiT-based FLUX.1. More results are available in the supplementary material.
@@ 4 Experiments
We validate the effectiveness of CoMPaSS on four popular open-weight text-to-image diffusion models. These models include three UNet-based architectures (SD1.4, SD1.5, and SD2.1) [58] and one MMDiT-based architecture (FLUX.1) [4]. Detailed experimental settings are provided in the supplementary material.
4.1 Evaluation Details
Table 2: Comparison of base models with their CoMPaSS-enhanced counterparts. Our method sets a new state of the art on all spatial benchmarks, including VISOR [26], T2I-CompBench Spatial [35] (T. Spat.), GenEval Position [25] (Pos.), and DPG-Bench Relation [33] (Rel.). We also report overall (Ovr.) scores for general compositional ability on GenEval and DPG-Bench, and image fidelity scores using FID [31] and CMMD [36]. Detailed per-benchmark results are in the supplementary material.
Method VISOR (%) T2I-CompBench GenEval DPG-Bench Image Fidelity
@@ 4.1 Evaluation Details
SD1.4 18.81 62.98 46.60 20.11 6.89 1.63 0.12 0.32 0.03 0.43 81.04 62.02 13.20 0.5618
SD1.4 +CoMPaSS 57.41 87.58 83.23 67.53 49.99 28.91 0.34 0.42 0.46 0.56 83.21 66.07 10.67 0.3313
SD1.5 17.58 61.08 43.65 18.62 6.49 1.57 0.08 0.31 0.04 0.42 73.49 63.18 12.82 0.5548
SD1.5 +CoMPaSS 61.46 93.43 86.55 72.13 54.64 32.54 0.35 0.42 0.54 0.57 84.10 65.81 10.89 0.3235
SD2.1 30.25 63.24 64.42 35.74 16.13 4.70 0.13 0.37 0.07 0.50 83.95 65.47 21.65 0.6472
SD2.1 +SPRIGHT [6] 43.23 71.24 71.78 51.88 33.09 16.15 0.21 - 0.11 0.51 - - 16.15 0.512
SD2.1 +CoMPaSS 62.06 90.96 85.02 71.29 56.03 35.90 0.32 0.43 0.51 0.54 86.54 69.48 16.96 0.4083
FLUX.1 37.96 66.81 64.00 44.18 28.66 14.98 0.18 0.46 0.26 0.60 92.30 80.63 27.96 0.8737
FLUX.1 +CoMPaSS 75.17 93.22 91.73 83.31 72.21 53.41 0.30 0.55 0.60 0.76 94.12 84.42 26.40 0.6859
Table 3: Computational efficiency comparison. Overheads exclude bounding box generation for methods that require it, and are reported in the format “Time/VRAM”. Lower is better.
Method Training Inference
SPRIGHT [6] +8.1%/+24.5% -
R&B [79] - +1108.7%/+150.9%
Attn. Refocusing [52] - +23.8%/+38.0%
Layout Guidance [13] - +180.9%/+67.4%
CoMPaSS (ours) +3.9%/+0.7% +2.47%/+0.6%
@@
Method # Training Images G. Position
SD2.1 0 0.07
SD2.1 +SPRIGHT [6] $\sim$500 0.11
[…]
SD2.1 +CoMPaSS 28k 0.51
FLUX.1 0 0.26
[…]
FLUX.1 +CoMPaSS 28k 0.60
Baselines. We compare CoMPaSS against several baselines: 1) SPRIGHT [6] is a training-based method whose focus on spatial consistency is highly aligned with ours. 2) R&B [79], 3) Attention Refocusing [52], and 4) Layout Guidance [13] are inference-only methods for zero-shot layout control. These methods rely on additional image-space guidance, such as bounding boxes or masks, which are typically provided either through manual annotation or by external language models [23, 41, 15].
Metrics. We extensively evaluate CoMPaSS on several benchmarks with varying complexity: VISOR [26] is a specialized benchmark with 31,680 prompts focusing on spatial understanding. T2I-CompBench [35] and GenEval [25] are two benchmarks for evaluating open-world compositional text-to-image generation. DPG-Bench [33] is a comprehensive dataset consisting of 1,065 lengthy, dense prompts, designed to assess the intricate semantic alignment capabilities of text-to-image models. FID [31] and CLIP Maximum Mean Discrepancy (CMMD) [36] assess image fidelity, with FID focusing on photorealism and CMMD providing an alternative measure that better aligns with human preferences. Additionally, we measure time and VRAM overhead to evaluate the computational efficiency of each method.
@@ 4.2 Main Results
Spatial Understanding. We evaluate spatial relationship generation accuracy across multiple benchmarks and report them in Tab. 2. Adding CoMPaSS to existing diffusion models sets a new state of the art across all spatial-related metrics. Notably, when applied to the state-of-the-art open-weight model FLUX.1, CoMPaSS achieves substantial relative gains on VISOR (+98%), T2I-CompBench Spatial (+67%), GenEval Position (131%), and the best scores on the DPG-Bench Relation benchmark. As shown in Fig. 5, CoMPaSS effectively addresses these limitations, correctly rendering novel spatial configurations unseen during training and demonstrating a significant enhancement in spatial understanding.
General Generation Capability and Fidelity. In addition to spatial metrics, we evaluate performance on other compositional tasks from the T2I-CompBench [35], GenEval [25], and DPG-Bench [33] benchmarks. We also assess image fidelity using FID [31] and CMMD [36]. The overall results are reported in Tab. 2, with detailed per-task results provided in Tab. A11 in the supplementary material. While CoMPaSS is designed to target spatial performance, we find that it also improves overall alignment scores and image fidelity. We conjecture that in base models, spatial terms are often entangled with unrelated semantics due to flawed data. By disentangling these terms, CoMPaSS may also help the model better understand other aspects of language, resulting in these broader improvements. Importantly, on DPG-Bench, which contains 1,065 long natural language prompts, models enhanced with CoMPaSS achieve consistently higher overall scores, indicating that the enhanced spatial understanding generalizes well to longer and more complex text prompts.
Computational Efficiency. We compare the time and VRAM overhead of CoMPaSS to prior work during both training and inference, with results reported in Tab. 3. SPRIGHT [6], a prior training-based method focused on spatial consistency, finetunes the text encoder along with the diffusion network. However, this approach inevitably adds computational requirements (+8.1% in time and +24.5% in VRAM). Inference-only methods [79, 52, 13] typically enable gradient-based optimization at test time, incurring substantial overhead (e.g. time overhead ranges from 23.8% up to 1108.7%). In contrast, CoMPaSS only alters the cross-attention calculation and introduces no extra trainable parameters. It is therefore lightweight, incurring only $\sim$3% time overhead and less than 1% VRAM overhead during inference.
[…]
@@ 4.3 Ablation Studies
Table 5: Ablation study on the hyperparameters of SCOP. Our main experiments use the hyperparameter set $\{\tau_{\text{v}},\tau_{\text{u}},\tau_{\text{o}},\tau_{\text{s}}\}=\{0.2,2.0,0.3,0.5\}$, denoted as “Ours” in the table. The results show that performance is robust to variations in these hyperparameters.
Name $\tau_{\text{v}}$ $\tau_{\text{u}}$ $\tau_{\text{o}}$ $\tau_{\text{s}}$ Ours
Value 0.1 0.3 1.0 3.0 0.2 0.5 0.3 0.7
SD1.5 0.48 0.52 0.50 0.49 0.47 0.51 0.51 0.53 0.54
FLUX.1 0.55 0.56 0.59 0.54 0.54 0.58 0.58 0.59 0.60
Table 6: Ablation study on the components of CoMPaSS. (i) original models; (ii) trained with the SCOP dataset described in Fig. 3; (iii) our full method. T2I-CompBench Spatial (T. Spatial) and GenEval Position (G. Pos) scores are reported.
@@
Figure 6: Qualitative ablation study on the TENOR module. The TENOR module improves generalization to unseen prompts. (i) First row: original models. (ii) Second row: models trained only with the SCOP dataset. (iii) Third row: our full method. The left three columns show results from SD1.5; the right three columns show results from FLUX.1.
@@
To understand the contribution of each component, we conduct controlled experiments on two representative models: SD1.5, a popular UNet-based model, and FLUX.1-dev, a state-of-the-art MMDiT-based model.
Hyperparameters of SCOP. The SCOP data engine has four tunable hyperparameters. We empirically determine the optimal values to be $\{\tau_{\text{v}},\tau_{\text{u}},\tau_{\text{o}},\tau_{\text{s}}\}=\{0.2,2.0,0.3,0.5\}$ via grid search. In Tab. 5, we report the model’s sensitivity to each of these four hyperparameters by evaluating performance on nearby values. While our chosen hyperparameters yield optimal results, the model’s performance remains high across a range of nearby values, indicating the robustness of the SCOP engine.
@@
Effect of SCOP. In Tab. 6, comparing settings (i) and (ii) within each model reveals that SCOP alone contributes substantially to the spatial understanding of diffusion models. This finding supports our initial hypothesis that a primary cause of failure in spatial generation is the flawed nature of existing training data. This confirms that SCOP successfully curates high-quality data that directly improves the spatial understanding of diffusion models.
@@
Effect of TENOR. Tab. 6 shows that our full method (setting (iii) within each model) further improves the spatial accuracy. As shown in Fig. 6, we evaluate the models on text descriptions unseen during training across all three settings. While SCOP provides rich and accurate spatial priors, it is only after incorporating TENOR that the model generalizes better to challenging, unseen spatial configurations. This phenomenon connects closely to our initial finding about spatial understanding deficiency in text encoders (Tab. 1): when the semantic representations of completely opposite text descriptions are indistinguishable, the model is trained to align mixed image signals from seemingly similar text conditions, resulting in poor generalization beyond training data. TENOR addresses this issue by injecting token ordering information into the text-image attention operations. This ensures that structurally different prompts produce distinct conditioning signals, enabling the model to learn correct associations from the SCOP data and thereby achieve better zero-shot generation on unseen prompts.
5 Discussion and Conclusion
Figure 7: Results for “a tiny phone above a large couch”. Left: FLUX.1. Middle: FLUX.1+CoMPaSS. Right: Fine-tuning (100 steps, batch size 4) on FLUX.1+CoMPaSS with only size-contrast data filtered from adapted SCOP.
Limitations and Future Work. While CoMPaSS demonstrates robust performance, its understanding of spatial language is not exhaustive. Our framework primarily addresses ambiguity arising from noisy web data and encoder limitations, but does not yet model the rich, context-dependent nature of spatial language, a key topic in Qualitative Spatial Relations (QSR) [9, 8, 16]. For instance, rendering subjects with extreme size disparities remains a challenge (see Fig. 7, left and middle), as the interpretation of “above” or “beside” depends heavily on object context. Our preliminary experiments show that a targeted fine-tuning stage can mitigate this specific issue (see Fig. 7, right). Furthermore, our current approach is viewer-centric and does not handle object-centric frames (e.g., “to the child’s right hand side”). Fully integrating QSR principles and modeling both object-centric and context-dependent relations are important directions for future work.
Conclusion. We have presented CoMPaSS, a versatile framework that enhances spatial understanding in T2I models. CoMPaSS uses the SCOP engine to curate high-quality, spatially-unambiguous training data, and the TENOR module to preserve crucial token ordering information that text encoders often lose. Extensive experiments show that CoMPaSS sets a new state of the art on key spatial benchmarks, achieving substantial relative gains on VISOR (+98%), T2I-CompBench Spatial (+67%), and GenEval Position (+131%) without compromising general generation capabilities or image quality. By addressing a fundamental limitation in T2I models, CoMPaSS represents an important step towards reliable, spatially-aware image generation.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 62032011 and the State Key Laboratory of CAD&CG at Zhejiang University.